OpenStudy (anonymous):

Find the value of "r" so the line that passes through each pair of points has the given slope. #1. (-3,-4),(-5,R),M=-9/2 #2. (1,4),(r,5),m=undefined #3. (-5,r)(1,3),m=7/6

OpenStudy (lgbasallote):

#1 m means slope so... \[\large m = \frac{y_2 - y_1}{x_2 - x_1}\] \[\frac{1}{2} = \frac{7-r}{6- (-2)}\]

OpenStudy (lgbasallote):

got it??

OpenStudy (anonymous):

yes ok I got the equation. How do I solve for r?

OpenStudy (lgbasallote):

\[\frac{1}{2} = \frac{7-r}{6+2}\] \[\frac{1}{2} = \frac{7-r}{8}\] multiply 8 to both sides... \[\large \frac{8}{2} = 7-r\] \[4 = 7 - r\] \[r = 7 -4\] r = 3

OpenStudy (lgbasallote):

can you solve the other numbers?

OpenStudy (anonymous):

I do not understand. These problems. Look at this one. (-3,-4),(-5,r),m=-9/2 I attempted to figure this out and got r=-13 is this correct?

OpenStudy (anonymous):

I'm stuck on this one. (-5, r),(1,3),m=7/6

OpenStudy (ajprincess):

Use the formula @lgbasallote has given u.

OpenStudy (anonymous):

I attempted to use that formula in the other ones, however, I am still having a difficult time. I have revised the question with the ones I am still stuck on.

OpenStudy (ajprincess):

Can u tell me where u r stuck?

OpenStudy (anonymous):

Yes the 3 problems that I have placed above. I am currently working on (-5,r)(1,3),m=7/6

OpenStudy (ajprincess):